# Mapping the positions of Two-Level-Systems on the surface of a superconducting transmon qubit

### v1

@@ Acknowledgements
As an example of the resolution limit, Fig. S6 shows data of a TLS that is determined to reside near the qubit’s Josephson junctions. There is some uncertainty whether the TLS resides on the upper or lower branch of the DC-SQUID. While most observed tuning ratios indicate a position on the upper branch, the data for $\gamma_{\alpha}/\gamma_{\gamma}$ (lower left panel in Fig. Fig. S6) suggests a solution on the lower branch. This may be due to the fit error determining the relatively weak $\gamma$-gate response $\gamma_{\gamma}$. The SQUID branches are separated by a distance of 6 µm (see also Fig. 1d). We expect that in our experiment, the achieved resolution is of similar size, and believe that it could be reduced to a few micrometers by mentioned improvements in sample design and measurement protocol.
Fig. S6: Data of a TLS that is identified to reside on the qubit’s DC-SQUID, similar to Figs. 2b,c. a Differences between the measured TLS response strength ratio and the corresponding simulated E-field ratio (color-coded), plotted for four electrode combinations. b Difference sum $\sigma$ (Eq. 3, colorscale) over all 6 unique combinations of electrode pairs. The white circle marks the global minimum, placing the TLS on the upper branch of the DC-SQUID’s loop.

### v2

@@ Acknowledgements
As an example of the resolution limit, Fig. S8 shows data of a TLS that is determined to reside near the qubit’s Josephson junctions. There is some uncertainty whether the TLS resides on the upper or lower branch of the DC-SQUID. While most observed tuning ratios indicate a position on the upper branch, the data for $\gamma_{\alpha}/\gamma_{\gamma}$ (lower left panel in Fig. S8) suggests a solution on the lower branch. This may be due to the fit error determining the relatively weak $\gamma$-gate response $\gamma_{\gamma}$. The SQUID branches are separated by a distance of 6 µm (see also Fig. 1d). We expect that in our experiment, the achieved resolution is of similar size, and believe that it could be reduced to a few micrometers by mentioned improvements in sample design and measurement protocol.
We tested the impact of errors on the measured TLS-electrode coupling strengths $\gamma_{i}$ with Monte-Carlo-simulations. For this, a TLS position was manually selected, and the exact tuning ratios at this position were determined from the E-field simulation. To simulate measurement uncertainty, these exact ratios were multiplied by random factors that have a constant distribution in an interval of $1\pm\delta$, where the maximum variation $\delta$ ranges from 1% to 50%. From these noisy tuning ratios, the TLS position was then determined with the mapping algorithm.
For a TLS placed on the SQUID and for a TLS on the qubit island, Fig. S9 and Fig.S10 respectively show resulting TLS positions from 1000 iterations (panels a and d). The deviations from the exact TLS location are shown in panels b for different variation factors $\delta$, and in panel e for differently chosen cut-off of the minimum electric field $E_{\mathrm{min}}$ that limits the detection area. We find that the median deviation between estimated and exact TLS positions ranges between 2 µm and 8 µm for variation factors up to $\delta=50$%, while for factors of $\delta>20\%$, the TLS position on the SQUID can not anymore reliably distinguished from a position on the island or ground plane. When the detection area is not restricted at small values of $E_{\mathrm{min}}$, TLS are mostly placed further away from the electrodes (see Fig. S9d). Such solutions are invalid since at these positions, TLS would not be observed in our TLS spectroscopy, and because the electric fields from the four electrode are not necessarily parallel, e.g. in the gap between ground plane and qubit island.
Fig. S8: Data of a TLS that is identified to reside on the qubit’s DC-SQUID, similar to Figs. 2b,c. a b Difference sum $\sigma$ (Eq. 3) over all 6 unique combinations of electrode pairs. The white circle marks the global minimum, placing the TLS on the upper branch of the DC-SQUID’s loop.
Fig. S9: a Determined positions (green circles) of a TLS on the DC-SQUID, whose exact location is marked by the blue circle, for different levels of noise (variation factor $\delta$) in the determined tuning ratios. b Histograms of deviations from the exact TLS position for different levels of variation $\delta$. c Difference between simulated E-field strengths and TLS tunability as in Fig. 2c, where minima (blue pixels) indicate most likely TLS positions. d Same as a, but varying the value of $E_{\mathrm{min}}$ that limits the detection area. e Deviations from the exact TLS location vs. $E_{\mathrm{min}}$. For Figs. a, b, $E_{\mathrm{min}}=1$ V/m, and for Figs. d, e, the variation was fixed at $\delta=10\%$
Fig. S10: Determined positions (green circles) of a TLS on the qubit island (blue circle), similar to Fig. S9. a and d Positions for different variation factors $\delta$ and values of $E_{\mathrm{min}}$, respectively, b and e: Histograms of deviations from the exact TLS position for different levels of variation $\delta$, and for different values of $E_{\mathrm{min}}$, respectively. c Difference between simulated E-field strengths and TLS tunability.
Before analyzing the data, we excluded several detected traces of TLS whose coupling strengths to the upper electrodes ($\alpha$ and $\gamma$ in Fig. 1) were indistinguishable from zero so that their position can not be reliably determined with the current procedure. These traces stem from TLS that are located further away on the extended arm of the qubit island (see Fig. 1b). Accordingly, possible solutions for TLS positions were restricted to the $\times105 µm$-large window as shown in Fig. 1c.